# Learning Japanese with Jouzu: Interaction Outcomes with Stylized Dialogue Fictional Agents

### v1

@@ 4.1. User Background
In the pre-use survey, background information including native language, other languages spoken, age, and other language learning apps used was collected from users. Below, we present the background information of the users who completed the study and provided background information: 50 participants.
English was the most frequently reported native language, accounting for 19 responses, with bilingual variants such as Korean and English present. Other native languages included Chinese (17), Korean (3), Hindi (5), Tamil (1), urkish (1), Japanese (1), and Malayam (1), reflecting a linguistically diverse sample. The age distribution showed a concentration of younger participants, with an average age of 24.32 years (SD = 3.78). The age range extended from 20 to 43 years, with 50% of respondents aged 23 or younger and 75% under the age of 25, indicating that most participants were in their early twenties. 39 participants had previously used Duolingo. Other language learning apps used included Rosetta Stone (4), Busuu (3), Babbel (3), and Memrise (1).
@@ 4.2. Qualitative Analysis
Figure 2. Mean participant ratings by experience category with standard deviation error bars (Likert scale: 1–5).
Table 1. Group-level descriptive statistics of user ratings (Likert scale: 1–5)
Group Mean SD Min Max Avg Item SD N
Amount Learned 2.48 2.48 2.48 1.09 44
Characters 3.39 3.39 3.39 1.48 44
Comparison to Other Apps 3.10 0.46 2.51 3.74 1.12 214
Fun and Engagement 3.31 0.36 2.77 3.55 1.13 176
Usability 3.64 3.64 3.64 1.10 44
Visuals, Art, and Aesthetic 3.52 3.52 3.52 1.37 44
@@ 4.3.1. Native Mandarin User Ratings
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese (including Mandarin) and non-Chinese speakers. Across all six dimensions, Chinese speakers rated the app more favorably. Most notably, the Characters group showed a large gap (M = 4.29 vs. 2.97), suggesting strong engagement with stylized agents among Chinese users. Visuals, Art, and Aesthetic also received a much higher rating from Chinese speakers (M = 4.29 vs. 3.17), followed by higher scores for Fun and Engagement (M = 3.82 vs. 3.07). Usability and Comparison to Other Apps showed smaller but consistent differences, while the largest perceived gap in learning outcomes (M = 2.79 vs. 2.33) aligned with open-ended feedback. These results point to meaningful cultural or pedagogical variation in how users interact with character-driven language tools. See Table 2.
Figure 3 visualizes the mean rating differences between Chinese and non-Chinese participants across all experience groups, highlighting consistently higher ratings by Chinese users.
Figure 3. Mean group ratings for Chinese vs. non-Chinese native speakers.
@@
Table 3. Independent-sample t-tests by native language group
@@ 4.3.2. Differences by Japanese Proficiency Level
Figure 4. Mean group ratings by Japanese proficiency level.
@@
Figure 5. Heatmap of mean experience ratings across Japanese proficiency levels. Darker shades indicate higher user satisfaction per category.
@@ 5.2.1. Multiple Choice Performance Analysis
This analysis compares participant performance across two sets of graded Japanese language questions before and after app usage. Among participants who completed both quizzes, average scores decreased slightly from 6.50 to 4.76 out of 10. However, high-end performance remained stable: two participants achieved perfect scores in the post-test, and six scored eight or higher. This suggests that, while some users may have struggled with retention or quiz difficulty, others maintained or improved their performance. These results, combined with increased comfort shown in open-ended output, indicate that the app may have supported language exploration and engagement, even if structured gains varied.
@@
5.2.2. Open Paragraph
Users were instructed to optionally write a paragraph in Japanese about anything they wished to the best of their ability. Before using the app, five respondents of varying proficiency levels – beginner, intermediate, and native – wrote paragraphs in Japanese. For example, an intermediate speaker wrote the following:
{CJK}
UTF8min
@@ 5.2.2. Open Paragraph
In contrast, after usage of the app, nine students wrote a paragraph in Japanese. Of these nine students, three had previously self-identified as complete beginners with no exposure to Japanese.
@@
One participant who was already identified as a beginner with some experience did not attempt a paragraph before use, but constructed a comprehensive sentence after use. It is important to note that, while characters do not use polite Japanese, this student still wrote a sentence using polite Japanese mixed with Saya’s character language.
{CJK}
UTF8min
@@ 6. Results
Analysis of interaction logs further revealed differentiated usage strategies by proficiency level. Complete beginners commonly used English prompts to request translations or vocabulary (e.g., “What is ‘sword’ in Japanese?”), with agents responding in Japanese using character-specific phrasing. These users often relied on passive listening or hybrid input patterns. In contrast, intermediate and advanced users engaged in full-sentence Japanese conversations, using both text and speech input to simulate native dialogue interaction.
@@ 7. Discussion
While multiple choice test scores decreased slightly after app usage, this does not necessarily indicate a decline in learning outcomes. Jouzu emphasizes open-ended, affective interaction over rote memorization, and several users who showed weaker multiple choice performance nonetheless demonstrated stronger expressive ability in paragraph writing. Many participants who initially wrote little or no Japanese were later able to produce multi-sentence output using vocabulary, grammar, or stylistic features encountered in the app. This suggests that the app supports exploratory, confidence-building learning, which may not be fully captured by short-format assessments.
Overall, Jouzu’s design suggests that culturally stylized, emotionally expressive dialogue agents—when embedded in a multimodal learning environment—can drive engagement and support meaningful language practice. Future work should explore adaptive features to help beginner users benefit more fully from the system’s immersive elements.
@@ 8. Conclusion
Jouzu demonstrates that stylized, character-based interaction can enhance user engagement and cultural immersion in language learning, though structured learning support remains essential for beginner users. Future work will explore adaptive scaffolding for novice learners and investigate long-term impacts of character-driven dialogue on spoken fluency.
Acknowledgements
This research was initiated independently of the primary author’s academic responsibilities at Columbia University and was conducted without the use of university resources.
{CJK}
UTF8min

### v2

@@ 5.1. User Background
In the pre-use survey, background information including native language, other languages spoken, age, and other language learning apps used was collected from users. Below, we present the background information of the users who completed the study and provided background information: 50 participants. English was the most frequently reported native language, accounting for 19 responses, with bilingual variants such as Korean and English present. Other native languages included Chinese (17), Korean (3), Hindi (5), Tamil (1), Turkish (1), Japanese (1), and Malayalam (1), reflecting a linguistically diverse sample. The age distribution showed a concentration of younger participants, with an average age of 24.32 years (SD = 3.78). The age range extended from 20 to 43 years, with 50% of respondents aged 23 or younger and 75% under the age of 25, indicating that most participants were in their early twenties. 39 participants had previously used Duolingo. Other language learning apps used included Rosetta Stone (4), Busuu (3), Babbel (3), and Memrise (1).
@@ 5.2. Qualitative Analysis
These descriptives establish a baseline for our core question: how stylization relates to user experience in a high barrier learning context. Summarizing group means shows where the system does work and where friction remains, and it motivates the subgroup tests by language background and proficiency as well as our triangulation with qualitative accounts. The dimensions align with our constructs: usability as a prerequisite for continued use, visuals and characters as stylization, fun and engagement as motivation, and perceived learning as an outcome-adjacent perception that can diverge from engagement.
@@
Figure 3. Mean participant ratings by experience category with standard deviation error bars (Likert scale: 1–5).:
Table 1. Group-level descriptive statistics of user ratings (Likert scale: 1–5):
Group Mean SD Min Max Avg Item SD
Amount Learned 2.48 2.48 2.48 1.09
Characters 3.39 3.39 3.39 1.48
Comparison to Other Apps 3.10 0.46 2.51 3.74 1.12
Fun and Engagement 3.31 0.36 2.77 3.55 1.13
Usability 3.64 3.64 3.64 1.10
Visuals, Art, and Aesthetic 3.52 3.52 3.52 1.37
@@ 5.3.1. Native Mandarin User Ratings
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese (including Mandarin) and non-Chinese speakers. Across all six dimensions, Chinese speakers rated the app more favorably. Most notably, the Characters group showed a large gap (M = 4.29 vs. 2.97), suggesting strong engagement with stylized agents among Chinese users. Visuals, Art, and Aesthetic also received a much higher rating from Chinese speakers (M = 4.29 vs. 3.17), followed by higher scores for Fun and Engagement (M = 3.82 vs. 3.07). Usability and Comparison to Other Apps showed smaller but consistent differences, while the largest perceived gap in learning outcomes (M = 2.79 vs. 2.33) aligned with open-ended feedback. These results point to meaningful cultural or pedagogical variation in how users interact with character-driven language tools (see Table 2). Figure 4 visualizes the mean rating differences between Chinese and non-Chinese participants across all experience groups, highlighting consistently higher ratings by Chinese users.
Figure 4. Mean group ratings for Chinese vs. non-Chinese native speakers. The left bars (blue) represent native Chinese speakers, and the right bars (orange) represent non-Chinese native speakers.:
@@
Table 3. Independent-sample t-tests by native language group. CN = Chinese Language Native, NCN = Non-Chinese Language Native
@@ 5.3.2. Differences by Japanese Proficiency Level
Figure 5. Mean group ratings by Japanese proficiency level. From left to right, the bars represent complete beginner — i.e., the user has had no exposure to Japanese — (blue), beginner level (orange), intermediate level (green), and native or near-native level (red).:
@@
Figure 6. Heatmap of mean experience ratings across Japanese proficiency levels. Darker shades indicate higher user satisfaction per category.:
@@ 6.2.1. Multiple Choice Performance Analysis
Participants completed graded Japanese language questions before and after using the app. Prior to usage, only 16 participants attempted the quiz, with no perfect scores recorded. Following app usage, 34 participants completed the quiz – more than double the original number – and 2 participants (5.9%) achieved perfect scores. Additionally, the proportion of high scores (8–10) increased from 12.5% to 17.6%. While the proportion of lower scores (1–3) also rose from 31.3% to 41.2%, this probably reflects the participation of new learners who gained enough confidence to attempt the quiz but had not yet reached higher proficiency. These results suggest that Jouzu supported increased learner engagement and test participation, with gains evident at both the high and entry levels of performance.
@@
6.2.2. Open Paragraph Assessment
Participants demonstrated substantial improvements across all three rubric categories: vocabulary diversity, grammatical complexity, and stylistic/pragmatic appropriateness. Group means increased from M=1.75M{=}1.75 to M=3.25M{=}3.25 for both vocabulary and grammar, and from M=2.25M{=}2.25 to M=4.00M{=}4.00 for style. Average gains were +1.50+1.50 (vocabulary), +1.50+1.50 (grammar), and +1.75+1.75 (style). Standard deviations narrowed post-intervention (e.g., vocabulary $SD{=}1.98\rightarrow 1.28$), suggesting more consistent performance. Effect sizes were large in all cases (paired Cohen’s d=0.85d=0.851.061.06).
User-level improvements.
Five of eight users showed measurable improvements in all three categories (Users 1, 2, 4, 5, and 7). User 5 exhibited the largest overall gain, improving from a total of 0 to 13 points ($\Delta{=}+13$). User 7 improved from 0 to 9 ($\Delta{=}+9$); User 1 from 0 to 7 ($\Delta{=}+7$); User 2 from 9 to 15 ($\Delta{=}+6$); and User 4 from 0 to 3 ($\Delta{=}+3$). Users 3, 6, and 8 showed no changes (totals of 15, 11, and 11 respectively, before and after) See Table 6 for rubric scores.
Distribution and relationships.
At the category level, five participants improved in vocabulary, five in grammar, and five in style; no decreases were observed. Improvements were highly correlated (e.g., vocabulary–style r=0.94r{=}0.94; vocabulary–grammar r=0.80r{=}0.80; grammar–style r=0.92r{=}0.92). Lower starting scores predicted larger gains (e.g., vocabulary start vs. improvement r=0.76r{=}-0.76), suggesting that the intervention disproportionately benefited participants who began with weaker performance. Overall rubric sums increased from 46 to 84 ($\Delta{=}+38$). See Table 7 for per-user totals.
Table 6. Rubric scores before and after the intervention (N=8N{=}8). The labels are abbreviated as BM (before-intervention mean), AM (after-intervention mean), standard deviations (SD), IM (improvement mean), which is defined as (After-Before), and paired Cohen’s $d$.
Category BM (SD) AM (SD) IM Cohen’s $d$
Vocabulary Diversity 1.75 (1.98) 3.25 (1.28) +1.50 1.06
Grammatical Complexity 1.75 (1.98) 3.25 (1.67) +1.50 0.85
Stylistic / Pragmatic App. 2.25 (2.49) 4.00 (1.51) +1.75 0.96
Overall (sum) 46 84 +38
Table 7. Per-user totals (sum of the three categories) before and after the intervention, with deltas.
User Before Total After Total Delta
User 1 0 7 +7
User 2 9 15 +6
User 3 15 15 0
User 4 0 3 +3
User 5 0 13 +13
User 6 11 11 0
User 7 0 9 +9
User 8 11 11 0
Totals 46 84 +38
Paragraph Examination
In addition to quantitative scores, we examined the paragraphs produced by participants to better understand the nature of their progress. These qualitative samples illustrate differences in fluency, grammatical control, and stylistic choices across proficiency levels. For instance, an intermediate speaker wrote the following:
@@ 6.2.2. Open Paragraph Assessment
One participant who was already identified as a beginner with some experience did not attempt a paragraph before use but constructed a comprehensive sentence after use. It is important to note that, while characters do not use polite Japanese, this student still wrote a sentence using polite Japanese mixed with Saya’s character language.
@@ 7. Results
Analysis of interaction logs further revealed different usage strategies by proficiency level. Complete beginners commonly used English prompts to request translations or vocabulary (e.g., “What is ‘sword’ in Japanese?”), with agents responding in Japanese using character-specific phrasing. These users often relied on passive listening or hybrid input patterns. In contrast, intermediate and advanced users engaged in full-sentence Japanese conversations, using both text and speech input to simulate native dialogue interaction.
Participants completed graded Japanese language questions before and after using the app. Prior to usage, only 16 participants attempted the quiz, with no perfect scores recorded. After app usage, 34 participants – more than double the original number – completed the quiz, with 2 participants (5.9%) achieving perfect scores. The proportion of high scores (8–10) also increased from 12.5% to 17.6%, indicating vocabulary gains for some learners. While the percentage of lower scores (1–3) rose from 31.3% to 41.2%, this likely reflects new users gaining enough confidence to attempt the quiz despite not yet achieving mastery. These results suggest that Jouzu fostered both increased test participation and broader performance distribution, with gains at the upper end and expanded access at the entry level.
@@
In addition to these broad trends, our rubric analysis shows consistent patterns of improvement across participants. On average, scores increased by 1.5 points in vocabulary and grammar, and by 1.75 points in stylistic and pragmatic appropriateness, all with large effect sizes. Five of the eight participants improved in every category, with gains especially pronounced among those who began with lower initial scores. By contrast, participants who entered the study with higher baseline proficiency tended to maintain their levels rather than progress further. This suggests that the system was particularly effective in supporting learners who started with limited skills, while more advanced users primarily experienced consolidation rather than measurable growth.
@@ 8. Discussion
Multiple choice test results indicate both an increase in participation and signs of improvement at the top end of performance. After using the app, more users completed the quiz, and a greater proportion achieved high scores (8–10) or perfect results. While some users also scored in the lower range, this likely reflects the broader base of participants, including those who previously may not have attempted the test. These findings suggest that Jouzu encouraged learners across a range of proficiency levels to engage with structured assessment, and that some users experienced measurable vocabulary gains. Importantly, these outcomes complement the expressive growth observed in paragraph writing, reinforcing the system’s effectiveness in supporting both test-oriented recall and open-ended language use.
Jouzu emphasizes open-ended, affective interaction over rote memorization, and several users who showed weaker multiple choice performance nonetheless demonstrated stronger expressive ability in paragraph writing. Many participants who initially wrote little or no Japanese were later able to produce multi-sentence output using vocabulary, grammar, or stylistic features encountered in the app. This suggests that the app supports exploratory, confidence-building learning, which may not be fully captured by short-format assessments.
Overall, Jouzu’s design suggests that culturally stylized, emotionally expressive dialogue agents – when embedded in a multimodal learning environment – can drive engagement and support meaningful language practice. Future work should explore adaptive features to help beginner users benefit more fully from the system’s immersive elements.
8.1. Design Implications
The findings from this study suggest several concrete design directions for future conversational language learning agents. The incorporation of stylized agents can serve not just as a novelty, but as a mechanism for increasing user motivation and affective immersion, particularly among learners with intermediate or higher proficiency. These users often mirrored stylistic elements in their own output and remained engaged longer during conversational sessions. However, the same features that support expressive language play for advanced users can present barriers for novices. Accordingly, systems may benefit from adaptive scaffolding mechanisms that calibrate agent responses based on learner proficiency. For instance, stylization could be made toggleable or modified dynamically during the session to scaffold more accessible interaction while retaining affective richness.
  • Stylized dialogue agents can increase motivation and expressive risk-taking, particularly for intermediate learners.
  • Beginners need scaffolding to navigate character language; toggleable stylization or adaptive agent personas may improve accessibility.
  • Cultural alignment (e.g., anime style for East Asian learners) boosts affective engagement, but may limit generalizability.
8.2. Limitations
This study has several limitations that should be taken into account when interpreting results or attempting to generalize findings. First, the participant pool consisted of computer science students, many of whom already had familiarity with technology or anime aesthetics, which may not reflect broader learner populations. Second, participation in assessments was optional, and baseline proficiency data was self-reported, limiting control over baseline variation and motivation. Third, while stylization proved effective for our sample, it was tightly tied to Japanese media tropes; different cultural stylizations may be needed to engage other learner communities. The absence of a control group also limits causal claims about effectiveness relative to traditional methods. Finally, our evaluation of open-ended writing relied on a manually applied rubric rather than a standardized or automated metric, though we explicitly justified this approach for its categorical clarity.
@@ 9. Conclusion
Jouzu demonstrates that stylized, character-based interaction enhances engagement, confidence, and cultural immersion, and that this engagement aligns with measurable learning improvements. After use, participants showed increases at the upper end of multiple-choice performance and substantial gains in expressive writing (vocabulary diversity, grammatical complexity, and stylistic appropriateness), alongside broader participation in assessment. These results suggest that culturally resonant, affective design can transform engagement into practice that supports learning, while complementing, rather than replacing, structured instruction. Future work should add adaptive scaffolding for beginners and examine how sustained engagement with stylized agents impacts longer-term spoken fluency.
@@ C.2. Post-use Quiz
Appendix D Open-ended Paragraph Grading Rubric
The following rubric was applied for grading open-ended paragraph responses.
Analytic rubric for Japanese writing samples (1--5 scale)
Vocabulary Diversity
  1. 1:
    Very small pool of high-frequency words; heavy repetition.
  2. 3:
    Mix of common items with some topic-specific terms; moderate repetition and some paraphrase.
  3. 5:
    Broad range of lexical items including topic-specific and stylistic lexemes; paraphrase and word-family variation evident.
Grammatical Complexity
  1. 1:
    Isolated words/fragments or only simple SOV sentences; minimal particle variety.
  2. 3:
    Simple sentences plus occasional coordination/subordination (e.g., ~て, ~から); some clause chaining; emerging particle control.
  3. 5:
    Varied clause combining (relative/embedded clauses, conditionals, modality/aspect); consistent and context-appropriate particle usage.
Stylistic / Pragmatic Appropriateness
  1. 1:
    No discernible register/persona; inconsistent or inappropriate politeness; character-language markers absent.
  2. 3:
    Some stylistic markers (e.g., sentence-final particles, limited persona cues); partly consistent with task and audience.
  3. 5:
    Sustained register management and persona (e.g., pronouns, sentence-final forms, particles) appropriate to task/character language; stylistic choices enhance communicative intent.